OpenStudy (anonymous):

A student has the following grades: 80, 76, 78, 83, 83. Find the variance of the grades. 1) 2.8 2) 7.6 3) 4.2 4) 0

OpenStudy (anonymous):

\[mean = (80+76+78+83+83)/5\] \[mean = 80\] Now you find the variance: \[var = \sum((x_i-mean)^2)/N\] \[var = (0+16+4+9+9)/5\] \[var = 38/5 = 7.6\]